OpenStudy (anonymous):

8(x + 1) - 3(x + 4) = 7(2 - x)

OpenStudy (anonymous):

This simplifies to 8x+8 - 3x-12 = 14 - 7x Using the distributive property. Then, just add them together 5x-4=14-7x 12x=18 x=18/12 = 3.2

OpenStudy (anonymous):

Thank You i understand this much better now :)

OpenStudy (anonymous):

It was 3/2 sorry for that typo :D